### v1

# Subdifferentials and penalty approximations of the obstacle problem

Abstract
We consider a framework for approximating the obstacle problem through a penalty approach by nonlinear PDEs. By using tools from capacity theory, we show that derivatives of the solution maps of the penalised problems converge in the weak operator topology to an element of the strong-weak Bouligand subdifferential. We are able to treat smooth penalty terms as well as nonsmooth ones involving for example the positive part function $\max(0,\cdot)$. Our abstract framework applies to several specific choices of penalty functions which are omnipresent in the literature. We conclude with consequences to the theory of optimal control of the obstacle problem.
1 Introduction
A ubiquitous method to approximate solutions of the classical obstacle problem
$u\in H^{1}_{0}(\Omega),\;u\leq\psi:\langle-\Delta u-f,u-v\rangle\leq 0\qquad\forall v\in H^{1}_{0}(\Omega),\;v\leq\psi$ (1)
is by penalisation through a nonlinear PDE
$-\Delta u_{\rho}+\Lambda_{\rho}(u_{\rho}-\psi)=f_{\rho}.$ (2)
The equation (2) approximates the variational inequality (1) in the sense that its solutions satisfy $u_{\rho}\to u$ as $\rho\to 0$ provided $f_{\rho}\to f$. If we define the source-to-solution map $f_{\rho}\mapsto u_{\rho}$ of (2) by
$S_{\rho}\colon H^{-1}(\Omega)\to H_{0}^{1}(\Omega)$
and consider its derivative at $f_{\rho}$ in a direction $d$ denoted by $\alpha_{\rho}:=S_{\rho}^{\prime}(f_{\rho})(d)$, then we know that it satisfies the linearised equation
$-\Delta\alpha_{\rho}+\Lambda_{\rho}^{\prime}(u_{\rho}-\psi)(\alpha_{\rho})=d.$ (3)
A natural question arises concerning the convergence of the derivatives $\alpha_{\rho}=S_{\rho}^{\prime}(f_{\rho})(d)$ in the limit $\rho\to 0$, or more generally, the convergence of the operators $S_{\rho}^{\prime}(f_{\rho})\colon H^{-1}(\Omega)\to H^{1}_{0}(\Omega)$. The possibility that immediately comes to mind is that they converge to the directional derivative of the VI solution map $S\colon u\mapsto f$, i.e., to $S^{\prime}(f)$. If this were to be the case, the limit $\alpha(d)$ would satisfy
$\alpha(d)\in\mathcal{K}:\langle-\Delta\alpha(d)-d,\alpha(d)-v\rangle\leq 0\quad\forall v\in\mathcal{K}$
where $\mathcal{K}:=\{v\in H^{1}_{0}(\Omega):\text{$v\leq 0$ q.e. on $\{u=\psi\}$, $\langle-\Delta u-f,v\rangle=0$}\}$ is the critical cone at $u=S(f)$. In general, this cannot happen since $d\mapsto\alpha(d)$ is nonlinear (unless $\mathcal{K}$ is a subspace), whereas $d\mapsto\alpha_{\rho}(d)$ is often linear for all $\rho$ (e.g., if $\Lambda_{\rho}$ is Fréchet differentiable) and linearity would be preserved in the limit.
In this paper, we consider limits of the operator $S_{\rho}^{\prime}(f_{\rho})\colon d\mapsto\alpha_{\rho}$ for different choices of the penalty function $\Lambda_{\rho}$ and prove that (under weak conditions) they converge in the weak operator topology to elements of the strong-weak subdifferential of $S$, defined as
$\partial_{B}^{sw}S(f):=\left\{\begin{aligned} L\in\mathcal{L}(H^{-1}(\Omega),H^{1}_{0}(\Omega)):\;&\exists\{f_{n}\}\subset F_{S}:f_{n}\to f\in H^{-1}(\Omega),\\ &S^{\prime}(f_{n})\overset{\mathrm{WOT}}{\to}L\text{ in $\mathcal{L}(H^{-1}(\Omega),H^{1}_{0}(\Omega))$}\end{aligned}\right\},$
where $F_{S}$ is the set of all points in $H^{-1}(\Omega)$ at which $S$ is Gâteaux differentiable and $\overset{\mathrm{WOT}}{\to}$ refers to convergence in the weak operator topology, which we recall now.
Definition 1.1.
A sequence $\{L_{n}\}\subset\mathcal{L}(X,Y)$ of bounded linear operators between Banach spaces $X$ and $Y$ converges to a bounded linear operator $L\in\mathcal{L}(X,Y)$ in the weak operator topology if and only if $L_{n}x\rightharpoonup Lx$ in $Y$ for all $x\in X$. We write this as $L_{n}\overset{\mathrm{WOT}}{\to}L$.
Our method of proof relies on a recent characterisation of $\partial^{sw}_{B}S(f)$ from [RW] (see (37) below) involving so-called capacitary measures, which we will introduce in Section 4.2. In order to fulfil one of the conditions to utilise that characterisation (see also Remark 1.4), we assume the following.
Assumption 1.2 (Standing assumption on regularity).
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded open set in dimension $n\geq 2$. For the obstacle, we assume $\psi\in C(\bar{\Omega})\cap H^{1}(\Omega)$ with either $\psi\in H^{1}_{0}(\Omega)$ or $\psi>0$ on $\partial\Omega$.
Inspired by tradition and existing literature, we treat in particular the following specific examples [^1] We have chosen the superscripts $\mathrm{m}$ and $\mathrm{c}$ in (4) to stand for max and complementarity respectively; the reason for the former is clear and the latter is due to the fact that $\Lambda^{\mathrm{c}}_{\rho}$ is obtained from writing the VI as a complementarity system. The superscripts $\mathrm{sm}$ and $\mathrm{sc}$ in (5) are supposed to denote that these are smoothed versions of $\mathrm{m}$ and $\mathrm{c}$ respectively.:
$\Lambda_{\rho}^{\mathrm{m}}(u)=\frac{1}{\rho}u^{+}\qquad$ $\text{and}\qquad\Lambda_{\rho}^{\mathrm{c}}(u)=\frac{1}{\rho}(\rho\bar{\lambda}+u)^{+}$ (4)
where $\bar{\lambda}\in L^{\infty}(\Omega)$, $\bar{\lambda}\geq 0$ is given, and the corresponding smooth versions
$\Lambda_{\rho}^{\mathrm{sm}}(u)=\frac{1}{\rho}m_{\rho}(u)\qquad$ $\text{and}\qquad\Lambda_{\rho}^{\mathrm{sc}}(u)=\frac{1}{\rho}m_{\rho}(\rho\bar{\lambda}+u)$ (5)
where $m_{\rho}$ is a regularisation (satisfying 2.4) that smooths out the positive part function $(\cdot)^{+}$, see 2.7 for some concrete examples. In particular, we have in mind the commonly used regularisations from [MR2822818] and [KunischWachsmuthPathFollowing] respectively, see (13) and (14) for their definitions. We will denote solution maps as well as other maps that depend on the specific choice of $\Lambda$ with the superscript $\mathrm{m},\mathrm{c},\mathrm{sm}$ or $\mathrm{sc}$ as appropriate. Note that the choice $\bar{\lambda}\equiv 0$ yields $\Lambda_{\rho}^{\mathrm{m}}=\Lambda_{\rho}^{\mathrm{c}}$ and $\Lambda_{\rho}^{\mathrm{sm}}=\Lambda_{\rho}^{\mathrm{sc}}$. A typical choice of $\bar{\lambda}$ is $\bar{\lambda}:=(f+\Delta\psi)^{+}$, see [IKOCVIs, Theorem 3.2].
Our penalty term $\Lambda_{\rho}^{\mathrm{sc}}$ covers also the penalisation
$\Lambda_{\rho}^{\widetilde{\mathrm{sc}}}(u)=\tilde{m}_{\rho}(\bar{\lambda}+(1/\rho)u),$ (6)
(given a regularisation $\tilde{m}_{\rho}$) which is used frequently in the literature, see, e.g., [KunischWachsmuth, KunischWachsmuthPathFollowing, SchielaWachsmuth, IK, IKOCVIs]. Indeed, $\Lambda_{\rho}^{\widetilde{\mathrm{sc}}}$ is of the form (5) using
$m_{\rho}(r):=\rho\tilde{m}_{\rho}(r/\rho),$ (7)
see Remark 2.6 for more details.
Let us give the main results of this work. We start with the smooth case (5). Below, we use the notation $C_{0}(\Omega)$ for the set of functions $v\in C(\bar{\Omega})$ with $v=0$ on $\partial\Omega$.
Theorem 1.3.
For every $f\in H^{-1}(\Omega)$ with $S(f)\in C_{0}(\Omega)$, if $f_{\rho}\to f$ in $H^{-1}(\Omega)$, then there exist maps $L^{\mathrm{sm}},L^{\mathrm{sc}}\in\partial_{B}^{sw}S(f)$ such that for a subsequence (that we relabel),
$(S_{\rho}^{\mathrm{sm}})^{\prime}(f_{\rho})\overset{\mathrm{WOT}}{\to}L^{\mathrm{sm}},\qquad\text{and}\qquad(S_{\rho}^{\mathrm{sc}})^{\prime}(f_{\rho})\overset{\mathrm{WOT}}{\to}L^{\mathrm{sc}}.$
Theorem 1.3 is a special case of Theorem 4.6 below.
Remark 1.4 (On the assumption $S(f)\in C_{0}(\Omega)$).
Note that asking for the solution $u=S(f)$ of (1) to satisfy $u\in C_{0}(\Omega)$ (which is needed for the characterisation of [RW, Lemma 4.3]) is not too restrictive. For example, when $\Omega$ is Lipschitz, [Rodrigues, Theorem 2.7, §5] guarantees $u\in C(\bar{\Omega})\cap H^{1}_{0}(\Omega)$ if $\psi\in C(\bar{\Omega})$ and $f\in W^{-1,p}(\Omega)$ for $p>n$. An alternative is if $\Omega$ satisfies a uniform exterior cone condition and $n\leq 3$, $\psi\in H^{1}(\Omega)$ with $\Delta\psi\in L^{2}(\Omega)$ and $f\in L^{2}(\Omega)$: then $u$ is even Hölder continuous, see [NeitzelWachsmuth, Theorem 2.7].
In the nonsmooth case (4), we additionally have to assume Gâteaux differentiability of the approximations.
Theorem 1.5.
For every sequence $\{f_{\rho}\}\subset H^{-1}(\Omega)$ such that $S_{\rho}^{\mathrm{m}}$ ($S_{\rho}^{\mathrm{c}}$) is Gâteaux differentiable at $f_{\rho}$ and $f_{\rho}\to f$ with $S(f)\in C_{0}(\Omega)$, there exists a map $L^{\mathrm{m}}\in\partial_{B}^{sw}S(f)$ ($L^{\mathrm{c}}\in\partial_{B}^{sw}S(f)$) such that for a subsequence (that we relabel),
$(S_{\rho}^{\mathrm{m}})^{\prime}(f_{\rho})\overset{\mathrm{WOT}}{\to}L^{\mathrm{m}},\qquad((S_{\rho}^{\mathrm{c}})^{\prime}(f_{\rho})\overset{\mathrm{WOT}}{\to}L^{\mathrm{c}}).$
Theorem 1.5 is also a special case of Theorem 4.6 below.
The nonsmooth result Theorem 1.5 contains the assumption that each $S_{\rho}$ is Gâteaux differentiable at $f_{\rho}$, which can be justified in the following sense. We rely on the key observation that there exists a dense set $F\subset H^{-1}(\Omega)$ such that $S_{\rho}^{\mathrm{m}}$ and $S_{\rho}^{\mathrm{c}}$ are Gâteaux differentiable from $F$ into $H^{1}_{0}(\Omega)$ for every $\rho$ taken from a countable set (see 4.2). Then the above theorem can always be applied for every constant sequence $f_{\rho_{n}}\equiv f\in F$ such that $S(f)\in C_{0}(\Omega)$.
In order to cover these cases without repetition and to generalise the structure of the penalty term as much as possible, we consider an abstract problem formulation, as described in the next subsection. In fact, as mentioned, Theorem 1.3 and Theorem 1.5 are consequences of our more general result Theorem 4.6. Finally, we also mention Theorem 5.1 where we obtain a first-order stationarity condition for an optimal control problem with a VI constraint.
Remark 1.6 (Generalisation to other elliptic operators).
In this paper we consider the elliptic operator in (1) and related problems to be the Laplacian because the characterisation of $\partial_{B}^{sw}S(f)$ from [RW] was shown for the Laplacian. Our results up to (but not including) the proof of Theorem 4.6 work in a more general setting where $-\Delta$ is replaced by a linear, bounded, coercive and T-monotone operator $A\colon H^{1}_{0}(\Omega)\to H^{-1}(\Omega)$. The estimates below should be adjusted to include the coercivity and boundedness constants.
2 Abstract setup and properties of the penalised problem
Throughout, we equip the Sobolev space $H_{0}^{1}(\Omega)$ with the inner product
$(u,v)_{H_{0}^{1}(\Omega)}:=\int_{\Omega}\nabla u\cdot\nabla v\mathop{}\!\mathrm{d}x.$
2.1 Setup
For the results of this section, it suffices to take an obstacle $\psi\in H^{1}(\Omega)$ with $\psi|_{\partial\Omega}\geq 0$ in the sense that $\min(0,\psi)\in H_{0}^{1}(\Omega)$.
Let us now formulate an abstract penalty term. For each $\rho>0$, we work with a general mapping $\Lambda_{\rho}\colon H^{1}_{0}(\Omega)\to H^{-1}(\Omega)$ defined as a Nemytskii map of a function $\lambda_{\rho}\colon\Omega\times\mathbb{R}\to\mathbb{R}$, i.e.,
$\Lambda_{\rho}(u)(x):=\lambda_{\rho}(x,u(x)).$
We make the following standing assumption on $\lambda_{\rho}$.
Assumption 2.1.
We assume that
  1. (i)
    for all $\rho\in(0,\infty)$, $\lambda_{\rho}\colon\Omega\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function,
  2. (ii)
    for all $\rho\in(0,\infty)$, $\lambda_{\rho}(x,\cdot)\colon\mathbb{R}\to\mathbb{R}$ is increasing and convex for a.a. $x\in\Omega$,
  3. (iii)
    for all $\rho\in(0,\infty)$, there exist $k_{\rho},K_{\rho}\in L^{\infty}(\Omega)$ and $j_{\rho}\in L^{2}(\Omega)$ with $k_{\rho}\leq K_{\rho}$ such that
    $\lambda_{\rho}(x,r)=\begin{cases}0&\text{if }r\leq k_{\rho}(x),\\ \frac{r+j_{\rho}(x)}{\rho}&\text{if }r\geq K_{\rho}(x),\end{cases}$ (8)
    and, for some $C>0$,
    $k_{\rho}$ $\to 0\text{ in }L^{\infty}(\Omega)\text{ as }\rho\to 0,$ (9)
    $\left\lVert k_{\rho}\right\rVert_{L^{2}(\Omega)}+\left\lVert K_{\rho}\right\rVert_{L^{2}(\Omega)}$ $\leq C\rho\quad\forall\rho\in(0,1],$ (10)
  4. (iv)
    $\lambda_{\rho}(x,\cdot)\in C^{1}(\mathbb{R})$ for all $\rho\in(0,\infty)$ and a.a. $x\in\Omega$ or $k_{\rho}\equiv K_{\rho}$.
Throughout this work, bearing in mind 2.1 (iv), we refer to the case that $\lambda_{\rho}(x,\cdot)\in C^{1}(\mathbb{R})$ for a.a. $x\in\Omega$ as the “smooth case” and to $k_{\rho}\equiv K_{\rho}$ as the “nonsmooth case”.
Remark 2.2.
  1. (i)
    The smoothness $\lambda_{\rho}(x,\cdot)\in C^{1}(\mathbb{R})$ is equivalent to $k_{\rho}(x)<K_{\rho}(x)$.
  2. (ii)
    Since we know that $\lambda_{\rho}(x,\cdot)\geq 0$, we have $\lambda_{\rho}(x,K_{\rho}(x))\geq 0$, which yields
    $K_{\rho}+j_{\rho}\geq 0.$ (11)
Lemma 2.3.
2.1 is satisfied for the nonsmooth choices in (4).
Proof.
Observe that for $\lambda_{\rho}^{\mathrm{m}}(r)=\frac{1}{\rho}r^{+}$,
$k_{\rho}=K_{\rho}=j_{\rho}=0,$
whereas for $\lambda_{\rho}^{\mathrm{c}}(x,r)=(1/\rho)(\rho\bar{\lambda}(x)+r)^{+}$, we have
$k_{\rho}(x)=K_{\rho}(x)=-\rho\bar{\lambda}(x),\quad j_{\rho}(x)=\rho\bar{\lambda}(x).$
From this information we can verify the claim without difficulty. ∎
Regarding the smooth cases, let us first give sufficient conditions on the structure of $m_{\rho}$ that will enable us to verify 2.1 more easily.
Assumption 2.4.
Let $m_{\rho}\in C^{1}(\mathbb{R})$ be given for all $\rho>0$, such that there exist $\theta_{\rho},\Theta_{\rho},l_{\rho}\in\mathbb{R}$ with
$m_{\rho}(r)=0\quad\text{for all }r\leq\theta_{\rho},\qquad m_{\rho}(r)=r+l_{\rho}\quad\text{for all }r\geq\Theta_{\rho},$
for all $\rho>0$ and
$|\theta_{\rho}|+|\Theta_{\rho}|\leq C\rho\quad\forall\rho\in(0,1]$ (12)
for some constant $C>0$.
Lemma 2.5.
If a function $m_{\rho}$ satisfies 2.4, then 2.1 is satisfied by both $\Lambda_{\rho}^{\mathrm{sm}}$ and $\Lambda_{\rho}^{\mathrm{sc}}$ defined as in (5).
Proof.
The first part of 2.4 directly implies that 2.1 (i), (ii) and (iv) are satisfied. The remaining part follows from observing that in the $\Lambda_{\rho}^{\mathrm{sm}}$ case, we have $k_{\rho}=\theta_{\rho}$, $K_{\rho}=\Theta_{\rho}$ and $j_{\rho}=l_{\rho}$, and in the $\Lambda_{\rho}^{\mathrm{sc}}$ case, we have $k_{\rho}=\theta_{\rho}-\rho\bar{\lambda}$, $K_{\rho}=\Theta_{\rho}-\rho\bar{\lambda}$ and $j_{\rho}=\rho\bar{\lambda}+l_{\rho}$.
Remark 2.6.
The penalisation $\Lambda_{\rho}^{\widetilde{\mathrm{sc}}}$ from (6) can be handled too and in fact we can in this situation weaken the last condition of 2.4: we would need the parameters of $\tilde{m}_{\rho}$ to satisfy only
$|\theta_{\rho}|+|\Theta_{\rho}|\leq C\quad\forall\rho\in(0,1].$
Indeed, if $\tilde{m}_{\rho}$ is a regularisation with the structure presented in 2.4 we find that $m_{\rho}$ defined as in (7) has associated parameters $\hat{\theta}_{\rho}:=\rho\theta_{\rho}$, $\hat{\Theta}_{\rho}:=\rho\Theta_{\rho}$ and $\hat{l}_{\rho}:=\rho l_{\rho}$, which all contain a helpful factor of $\rho$.
Let us now look at some examples of $m_{\rho}$ that satisfy 2.4.
Lemma 2.7.
The following choices of $m_{\rho}$ satisfy 2.4.
  1. (i)
    The global regularisation used, e.g., in [MR2822818]:
    $m_{\rho}(r):=\begin{cases}0&\text{if }r\leq 0,\\ \frac{r^{2}}{2\rho}&\text{if }0<r<\rho,\\ r-\frac{\rho}{2}&\text{if }r\geq\rho,\end{cases}$ (13)
    with $\theta_{\rho}=0$, $\Theta_{\rho}=\rho$, $l_{\rho}=-\rho/2$.
  2. (ii)
    The regularisation from [KunischWachsmuthPathFollowing]:
    $\tilde{m}_{\rho}(r):=\begin{cases}0&\text{if }r\leq-\frac{\rho}{2},\\ \frac{1}{2\rho^{3}}\left(r+\frac{\rho}{2}\right)^{3}\left(\frac{3\rho}{2}-r\right)&\text{if $-\frac{\rho}{2}<r<\frac{\rho}{2}$},\\ r&\text{if $r\geq\frac{\rho}{2}$},\end{cases}$ (14)
    with $\theta_{\rho}=-\frac{\rho}{2}$, $\Theta_{\rho}=\frac{\rho}{2}$, $l_{\rho}=0$.
  3. (iii)
    The local regularisation from [MR2822818]:
    $m_{\rho}^{l}(r):=\begin{cases}0&\text{if }r\leq-\rho,\\ \frac{r^{2}}{4\rho}+\frac{r}{2}+\frac{\rho}{4}&\text{if }-\rho<r<\rho,\\ r&\text{if }r\geq\rho,\end{cases}$
    with $\theta_{\rho}=-\rho$, $\Theta_{\rho}=\rho$, $l_{\rho}=0$.
  4. (iv)
    The regularisation from [KunischWachsmuth]:
    $\hat{m}_{\rho}(r):=\begin{cases}0&\text{if }r\leq-\frac{\rho}{2},\\ \frac{1}{2\rho}\left(r+\frac{\rho}{2}\right)^{2}&\text{if $-\frac{\rho}{2}<r<\frac{\rho}{2}$},\\ r&\text{if $r\geq\frac{\rho}{2}$},\end{cases}$
    with $\theta_{\rho}=-\frac{\rho}{2}$, $\Theta_{\rho}=\frac{\rho}{2}$, $l_{\rho}=0$.
Hence $\Lambda_{\rho}^{sm}$ and $\Lambda_{\rho}^{sc}$ as defined in (5) associated to each of the above regularisations satisfy 2.1.
Proof.
We can verify the claim using the information presented below each choice. ∎
2.2 First properties
Lemma 2.8.
For all $\rho\in(0,\infty)$ and a.a. $x\in\Omega$ the following holds.
  1. (i)
    The map $\lambda_{\rho}(x,\cdot)$ is directionally differentiable. In the nonsmooth case (with $k_{\rho}\equiv K_{\rho}$), the directional derivative is given by
    $\lambda_{\rho}^{\prime}(x,r)(h)=\frac{1}{\rho}\chi_{\{r=k_{\rho}\}}(x)h^{+}+\frac{1}{\rho}\chi_{\{r>k_{\rho}\}}(x)h\qquad\forall r,h\in\mathbb{R}.$ (15)
  2. (ii)
    In the smooth case we have
    $0\leq\lambda_{\rho}^{\prime}(x,r)\leq\frac{1}{\rho}\qquad\forall r\in\mathbb{R}.$ (16)
  3. (iii)
    We have
    $\lambda_{\rho}^{\prime}(x,k_{\rho}(x))(\alpha)\alpha\geq 0,\quad\lambda_{\rho}^{\prime}(x,k_{\rho}(x))(\alpha)\alpha^{+}$ $\geq 0$ $\forall\alpha\in\mathbb{R},$ (17)
    $|\lambda_{\rho}^{\prime}(x,r)(h)|$ $\leq\frac{1}{\rho}|h|$ $\forall r,h\in\mathbb{R}.$ (18)
  4. (iv)
    The function $\lambda_{\rho}(x,\cdot)$ is Lipschitz continuous uniformly in $x$, i.e.,
    $|\lambda_{\rho}(x,u)-\lambda_{\rho}(x,v)|\leq\frac{1}{\rho}|u-v|\qquad\forall u,v\in\mathbb{R}.$ (19)
  5. (v)
    If $\rho\in(0,1]$, we have the growth condition (with the constant $C$ from (10))
    $\left\lVert j_{\rho}\right\rVert_{L^{2}(\Omega)}\leq 2C\rho.$ (20)
  6. (vi)
    The map $\Lambda_{\rho}\colon L^{2}(\Omega)\to L^{2}(\Omega)$ is well defined, Lipschitz continuous and directionally differentiable with the derivative given by
    $\Lambda_{\rho}^{\prime}(u)(h)(x)=\lambda_{\rho}^{\prime}(x,u(x))(h(x)).$ (21)
  7. (vii)
    The map $\Lambda_{\rho}\colon L^{2}(\Omega)\to L^{2}(\Omega)$ is monotone.
  8. (viii)
    For all $r\in\mathbb{R}$ we have
    $\lambda_{\rho}(x,r)r^{+}\geq-\frac{K_{\rho}(x)}{\rho}r^{+}+\frac{1}{\rho}|r^{+}|^{2}.$ (22)
  9. (ix)
    If $\rho\in(0,1]$, we have (with the constant $C$ from (10))
    $\left\lVert\Lambda_{\rho}(v-\psi)\right\rVert_{L^{2}(\Omega)}\leq\frac{1}{\rho}\left\lVert(v-\psi)^{+}\right\rVert_{L^{2}(\Omega)}+C\qquad\forall v\in H_{0}^{1}(\Omega).$ (23)
Proof.
  1. (i)
    In the nonsmooth case, we have $k_{\rho}\equiv K_{\rho}$. Consequently, the formula for $\lambda_{\rho}(x,\cdot)$ yields the differentiability everywhere on points other than at $k_{\rho}$, where it is directionally differentiable; it is standard to see that (15) is the expression for the derivative.
  2. (ii)
    The fact that $\lambda_{\rho}(x,\cdot)$ is convex and differentiable implies that its derivative is increasing, so this follows from the structure of $\lambda_{\rho}$ given in 2.1 (iii).
  3. (iii)
    Regarding (17), for the nonnegativity, this is clear in the smooth case due to the linearity with respect to the direction and the nonnegativity of the derivative. In the nonsmooth case this follows from the expression (15) for the derivative above.
    For the upper bound (18), in the smooth case this follows again by (16). In the nonsmooth case this follows by the expression for the directional derivative above and the fact that the sets appearing in the expression are disjoint.
  4. (iv)
    Using the mean value theorem for directional derivatives [MR2986672, Proposition 2.29] and (18), we obtain the claim.
  5. (v)
    We note that by (19),
    $\frac{1}{\rho}|K_{\rho}+j_{\rho}|=|\lambda_{\rho}(\cdot,K_{\rho})-\lambda_{\rho}(\cdot,k_{\rho})|\leq\frac{1}{\rho}|K_{\rho}-k_{\rho}|$
    which implies $|j_{\rho}|\leq|K_{\rho}-k_{\rho}|+|k_{\rho}|$ whence (20) follows by (10).
  6. (vi)
    The above Lipschitz property implies
    $|\lambda_{\rho}(x,u)|=|\lambda_{\rho}(x,u)-\lambda_{\rho}(x,k_{\rho}(x))|\leq\frac{1}{\rho}|u-k_{\rho}(x)|$ (24)
    and thus $\Lambda_{\rho}$ maps $L^{2}(\Omega)$ to $L^{2}(\Omega)$. Lipschitz continuity also follows easily by the above.
    Due to the directional differentiability of $\lambda_{\rho}(x,\cdot)$ and a simple dominated convergence theorem argument, using the fact that the first derivative is bounded by (18), we obtain directional differentiability.
  7. (vii)
    This follows from the fact that $\lambda_{\rho}(x,\cdot)$ is increasing.
  8. (viii)
    From the basic properties of $\lambda_{\rho}$ and (11) we get
    $\lambda_{\rho}(x,r)\geq\frac{r-K_{\rho}(x)}{\rho}.$
    Indeed, for $r<K_{\rho}$, the right-hand side is negative and for $r\geq K_{\rho}$ this follows from (8) and (11). Now, (22) easily follows.
  9. (ix)
    Using the monotonicity of $\lambda_{\rho}(x,\cdot)$, $\lambda_{\rho}(x,k_{\rho}(x))=0$ and the Lipschitz estimate (19), we get
    $\lambda_{\rho}(x,v(x)-\psi(x))\leq\lambda_{\rho}(x,(v(x)-\psi(x))^{+})\leq\frac{1}{\rho}(v(x)-\psi(x))^{+}+\frac{1}{\rho}|k_{\rho}(x)|.$
    Taking the $L^{2}(\Omega)$-norm and using (10), the inequality (23) follows. ∎
Next, we address the well posedness of the PDE (2) satisfied by $u_{\rho}$.
Lemma 2.9.
For all $\rho>0$, the solution map $S_{\rho}\colon H^{-1}(\Omega)\to H^{1}_{0}(\Omega)$ of the PDE (2) is well defined. If $\rho\in(0,1]$, we have
$\left\lVert S_{\rho}(f_{\rho})\right\rVert_{H_{0}^{1}(\Omega)}$ $\leq\left\lVert f_{\rho}\right\rVert_{H^{-1}(\Omega)}+2\left\lVert\min(0,\psi)\right\rVert_{H_{0}^{1}(\Omega)}+C_{P}C\qquad\forall f_{\rho}\in H^{-1}(\Omega),$
where $C_{P}$ is the constant from Poincaré’s inequality and $C$ is from (10). Furthermore, $S_{\rho}\colon H^{-1}(\Omega)\to H^{1}_{0}(\Omega)$ is Lipschitz continuous with constant $1$.
Proof.
We define the operator $T_{\rho}\colon H_{0}^{1}(\Omega)\to H_{0}^{1}(\Omega)$ via
$T_{\rho}(u):=(-\Delta)^{-1}\Lambda_{\rho}(u-\psi)\qquad\forall u\in H_{0}^{1}(\Omega).$
By applying $(-\Delta)^{-1}\colon H^{-1}(\Omega)\to H_{0}^{1}(\Omega)$ to (2), we get
$(\operatorname{Id}+T_{\rho})(u_{\rho})=(-\Delta)^{-1}f_{\rho}.$
Now, it is easy to check that the operator $T_{\rho}$ is continuous and monotone. Consequently, the operator $T_{\rho}$ is maximally monotone, see [BauschkeCombettes2011, Proposition 20.27]. Thus, we can apply Minty’s theorem, see [BauschkeCombettes2011, Theorem 21.1 and Proposition 23.8], to obtain that $\operatorname{Id}+T_{\rho}$ bijective and the inverse has Lipschitz constant $1$. Since $-\Delta\colon H_{0}^{1}(\Omega)\to H^{-1}(\Omega)$ is an isometric isomorphism, this shows that $S_{\rho}$ is well defined and has Lipschitz constant $1$.
It remains to provide an estimate for $S_{\rho}(f_{\rho})$. We define $\psi_{0}:=\min(0,\psi)$ and recall that $\psi_{0}\in H_{0}^{1}(\Omega)$ by our assumption on $\psi$. We observe $S_{\rho}(g_{\rho})=\psi_{0}$ for $g_{\rho}:=-\Delta\psi_{0}+\Lambda_{\rho}(\psi_{0}-\psi)$. Consequently, the Lipschitzness of $S_{\rho}$ yields
$\left\lVert S_{\rho}(f_{\rho})\right\rVert_{H_{0}^{1}(\Omega)}$ $\leq\left\lVert S_{\rho}(f_{\rho})-S_{\rho}(g_{\rho})\right\rVert_{H_{0}^{1}(\Omega)}+\left\lVert S_{\rho}(g_{\rho})\right\rVert_{H_{0}^{1}(\Omega)}$
$\leq\left\lVert f_{\rho}+\Delta\psi_{0}-\Lambda_{\rho}(\psi_{0}-\psi)\right\rVert_{H^{-1}(\Omega)}+\left\lVert\psi_{0}\right\rVert_{H^{1}_{0}(\Omega)}$
$\leq\left\lVert f_{\rho}\right\rVert_{H^{-1}(\Omega)}+2\left\lVert\psi_{0}\right\rVert_{H_{0}^{1}(\Omega)}+C_{P}\left\lVert\Lambda_{\rho}(\psi_{0}-\psi)\right\rVert_{L^{2}(\Omega)}.$
The last addend is estimated via (23) by using $(\psi_{0}-\psi)^{+}=0$. ∎
Let us now address the differentiability of $S_{\rho}$.
Lemma 2.10.
Given $f_{\rho},d\in H^{-1}(\Omega)$, the directional derivative $\alpha_{\rho}=S_{\rho}^{\prime}(f_{\rho})(d)$ exists and is the unique solution of the PDE (3), i.e.,
$-\Delta\alpha_{\rho}+\Lambda_{\rho}^{\prime}(u_{\rho}-\psi)(\alpha_{\rho})=d$
and satisfies the bound
$\left\lVert\alpha_{\rho}\right\rVert_{H^{1}_{0}(\Omega)}\leq\left\lVert d\right\rVert_{H^{-1}(\Omega)}.$
Proof.
The directional differentiability and satisfaction of (3) follows by the directional differentiability and monotonicity of $\Lambda_{\rho}$, see e.g. [AHRW, Lemma 6.1] or [CCMW, Theorem 2.2].
For the estimate, we test (3) with $\alpha_{\rho}$ and obtain
$\left\lVert\alpha_{\rho}\right\rVert_{H^{1}_{0}(\Omega)}^{2}+\int_{\Omega}\lambda_{\rho}^{\prime}(\cdot,u_{\rho}-\psi)(\alpha_{\rho})\alpha_{\rho}\mathop{}\!\mathrm{d}x=\langle d,\alpha_{\rho}\rangle.$
Using the nonnegativity by (17) we get the result.
Uniqueness follows by the monotonicity of $\Lambda_{\rho}^{\prime}(u_{\rho}-\psi)\colon H^{1}_{0}(\Omega)\to H^{-1}(\Omega)$, which is a consequence of the fact that the derivative derivative of a monotone map at a particular point is also monotone. ∎
In what follows, we show that $\Lambda_{\rho}^{\prime}(u_{\rho}-\psi)$ can be represented by a function in $L^{\infty}(\Omega)$ under the assumption that the directional derivative $S_{\rho}^{\prime}(f_{\rho})$ is linear.
Lemma 2.11.
If $f_{\rho}\in H^{-1}(\Omega)$ is given such that $S_{\rho}^{\prime}(f_{\rho})$ is linear, i.e., $S_{\rho}^{\prime}(f_{\rho})\in\mathcal{L}(H^{-1}(\Omega),H^{1}_{0}(\Omega))$, then for a.a. $x\in\Omega$, $\lambda_{\rho}^{\prime}(x,\cdot)$ is differentiable at the point $S_{\rho}(f_{\rho})(x)-\psi(x)$ and the derivative belongs to $[0,1/\rho]$. Consequently, the operator $\Lambda_{\rho}^{\prime}(S_{\rho}(f_{\rho})-\psi)\colon L^{2}(\Omega)\to L^{2}(\Omega)$ is linear and can be identified with a function in $L^{\infty}(\Omega)$.
Proof.
The operator
$S_{\rho}^{\prime}(f_{\rho})\colon H^{-1}(\Omega)\to H_{0}^{1}(\Omega)$
is the inverse of
$-\Delta+\Lambda_{\rho}^{\prime}(S_{\rho}(f_{\rho})-\psi)\colon H_{0}^{1}(\Omega)\to H^{-1}(\Omega).$
Since the former operator is linear by assumption, the latter operator is linear as well. Consequently, $\Lambda_{\rho}^{\prime}(S_{\rho}(f_{\rho})-\psi)$ is a linear operator. From (21) we infer the differentiability of $\lambda_{\rho}^{\prime}(x,\cdot)$ at $S_{\rho}(f_{\rho})(x)-\psi(x)$ and then (15), (16) imply $\lambda_{\rho}^{\prime}(\cdot,S_{\rho}(f_{\rho})-\psi)\in L^{\infty}(\Omega)$. ∎
In this linear case, we can define a measure $\mu_{\rho}$ via
$\mu_{\rho}(B):=\int_{B}\Lambda_{\rho}^{\prime}(S_{\rho}(f_{\rho})-\psi)\mathop{}\!\mathrm{d}x$ (25)
and (3) becomes
$-\Delta\alpha_{\rho}+\alpha_{\rho}\mu_{\rho}=d,$ (26)
with an appropriate interpretation of the term $\alpha_{\rho}\mu_{\rho}$. We will use the notion of capacitary measures to pass to the limit in this equation, see Section 4.2.
2.3 Convergence results
We now consider the limit $\rho\to 0$. The convergence result
$u_{\rho}\to u:=S(f)\qquad\text{in }H_{0}^{1}(\Omega)$
has been shown for various choices of the penalty function in, e.g., [AHRW, Lemma 3.3 and Lemma 3.5], [IK, Theorem 3.1], [IKOCVIs, Theorem 4.1], [IKH1, Theorem 2.1] and [SchielaWachsmuth, Theorem 2.10]. We demonstrate that this convergence already holds under our general setting. In particular, our results are more general than currently available in the literature and are phrased in the natural function spaces for the problem (e.g., the source terms are only assumed to converge in the dual space). We begin with a preparatory lemma.
Lemma 2.12.
For all $f_{\rho}\in H^{-1}(\Omega)$ and $\rho\in(0,1]$, we have
$\left\lVert(S_{\rho}(f_{\rho})-\psi)^{+}\right\rVert_{L^{2}(\Omega)}$ $\leq\frac{\sqrt{\rho}}{2}\left(\left\lVert f_{\rho}+\Delta\psi\right\rVert_{H^{-1}(\Omega)}+C_{P}C\right),$ (27)
$\left\lVert\Lambda_{\rho}(S_{\rho}(f_{\rho})-\psi)\right\rVert_{L^{2}(\Omega)}$ $\leq\frac{1}{2\sqrt{\rho}}\left(\left\lVert f_{\rho}+\Delta\psi\right\rVert_{H^{-1}(\Omega)}+C_{P}C\right)+C,$ (28)
where $C$ is the constant from (10).
Proof.
Denoting $u_{\rho}:=S_{\rho}(f_{\rho})$, we test the equation
$-\Delta(u_{\rho}-\psi)+\Lambda_{\rho}(u_{\rho}-\psi)=f_{\rho}+\Delta\psi$
with $(u_{\rho}-\psi)^{+}\in H^{1}_{0}(\Omega)$. One has (by T-monotonicity of the Laplacian)
$\langle-\Delta(u_{\rho}-\psi),(u_{\rho}-\psi)^{+}\rangle\geq\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{H^{1}(\Omega)}^{2}$
and using (22) we obtain
$\int_{\Omega}\Lambda_{\rho}(u_{\rho}-\psi)(u_{\rho}-\psi)^{+}\mathop{}\!\mathrm{d}x\geq\frac{1}{\rho}\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{L^{2}(\Omega)}^{2}-\frac{1}{\rho}\int_{\Omega}K_{\rho}(u_{\rho}-\psi)^{+}\mathop{}\!\mathrm{d}x,$
which leads to
$\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{H^{1}_{0}(\Omega)}^{2}+\frac{1}{\rho}\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{L^{2}(\Omega)}^{2}\leq\Big{\langle}f_{\rho}+\Delta\psi+\frac{K_{\rho}}{\rho},(u_{\rho}-\psi)^{+}\Big{\rangle}$ (29)
$\leq\frac{1}{4}\left\lVert f_{\rho}+\Delta\psi+\frac{K_{\rho}}{\rho}\right\rVert_{H^{-1}(\Omega)}^{2}+\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{H_{0}^{1}(\Omega)}^{2},$
where we used Young’s inequality. This yields
$\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{L^{2}(\Omega)}\leq\frac{\sqrt{\rho}}{2}\left\lVert f_{\rho}+\Delta\psi+\frac{K_{\rho}}{\rho}\right\rVert_{H^{-1}(\Omega)}$
and (27) follows from Poincaré’s inequality. Inequality (28) follows with (23). ∎
Remark 2.13.
By arguing as in the proof of [SchielaWachsmuth, Lemma 2.3] we can obtain a better rate than above if we have additional regularity. Indeed, we can show that if $f_{\rho}+\Delta\psi+K_{\rho}/\rho\in L^{2}(\Omega)$ and $\rho\in(0,1]$, we have that
$\left\lVert(S_{\rho}(f_{\rho})-\psi)^{+}\right\rVert_{L^{2}(\Omega)}$ $\leq\rho\left\lVert f_{\rho}+\Delta\psi+\frac{K_{\rho}}{\rho}\right\rVert_{L^{2}(\Omega)}$ (30)
$\left\lVert\Lambda_{\rho}(S_{\rho}(f_{\rho})-\psi)\right\rVert_{L^{2}(\Omega)}$ $\leq\left\lVert f_{\rho}+\Delta\psi+\frac{K_{\rho}}{\rho}\right\rVert_{L^{2}(\Omega)}+C$ (31)
where $C$ is the constant from (10). To see this, we can adapt (29) as follows, making use of the assumed $L^{2}$ regularity:
$\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{H^{1}_{0}(\Omega)}^{2}+\frac{1}{\rho}\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{L^{2}(\Omega)}^{2}\leq\Big{\langle}f_{\rho}+\Delta\psi+\frac{K_{\rho}}{\rho},(u_{\rho}-\psi)^{+}\Big{\rangle}$
$\leq\left\lVert f_{\rho}+\Delta\psi+\frac{K_{\rho}}{\rho}\right\rVert_{L^{2}(\Omega)}\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{L^{2}(\Omega)},$
whence (30). The second estimate (31) follows as before.
Proposition 2.14.
Let $f_{\rho}\to f$ in $H^{-1}(\Omega)$. We have $u_{\rho}=S_{\rho}(f_{\rho})\to S(f)=:u$ in $H^{1}_{0}(\Omega)$, i.e., the limit $u$ solves the VI (1).
Proof.
The estimate in 2.9 implies that $\{u_{\rho}\}$ is uniformly bounded in $H^{1}_{0}(\Omega)$, hence there exists a $v\in H^{1}_{0}(\Omega)$ such that for a subsequence (that we relabel), we have
$u_{\rho}\rightharpoonup v\quad\text{in $H^{1}_{0}(\Omega)$}.$
In what follows, we argue that the convergence is strong and that the limit $v$ equals $u$. Hence, the entire sequence converges towards $u:=S(f)$ in $H_{0}^{1}(\Omega)$.
We first note that (27) implies the feasibility of $v$, i.e., $v\leq\psi$. Consequently, we can use $v$ in the VI (1) and test the PDE for $u_{\rho}$ with $u_{\rho}-u$. Adding the resulting expressions gives
$\left\lVert u-u_{\rho}\right\rVert_{H_{0}^{1}(\Omega)}^{2}\leq\langle f-f_{\rho},u-u_{\rho}\rangle+\langle-\Delta u-f,v-u_{\rho}\rangle-\langle\Lambda_{\rho}(u_{\rho}-\psi),u_{\rho}-u\rangle.$
It remains to check that the last addend on the right-hand side converges to 00. From $v-\psi+k_{\rho}\leq k_{\rho}$ we get $\Lambda_{\rho}(v-\psi+k_{\rho})=0$, see (8). Together with the monotonicity of $\Lambda_{\rho}$, we obtain
$\langle\Lambda_{\rho}(u_{\rho}-\psi),u_{\rho}-v\rangle$ $=\langle\Lambda_{\rho}(u_{\rho}-\psi)-\Lambda_{\rho}(v-\psi+k_{\rho}),u_{\rho}-(v+k_{\rho})+k_{\rho}\rangle$
$\geq\langle\Lambda_{\rho}(u_{\rho}-\psi),k_{\rho}\rangle.$
Combining (28) with (10) yields that the right-hand side converges to 00. This can be used in the above estimate and the claim follows. ∎
Now, we can revisit the estimate in 2.12 and improve the rate implied by (27). This will be crucial later on.
Lemma 2.15.
Let $f_{\rho}\to f$ in $H^{-1}(\Omega)$. Then,
$\rho^{-1/2}\left\lVert(S_{\rho}(f_{\rho})-\psi)^{+}\right\rVert_{L^{2}(\Omega)}\to 0.$
Proof.
Thanks to 2.14, we have $(u_{\rho}-\psi)^{+}\to 0$ in $H_{0}^{1}(\Omega)$ with $u_{\rho}:=S_{\rho}(f_{\rho})$. In the inequality (29), i.e.,
$\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{H^{1}_{0}(\Omega)}^{2}+\frac{1}{\rho}\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{L^{2}(\Omega)}^{2}\leq\Big{\langle}f_{\rho}+\Delta\psi+\frac{K_{\rho}}{\rho},(u_{\rho}-\psi)^{+}\Big{\rangle},$
the duality product on the right-hand side clearly goes to zero, see (10). This shows the assertion. ∎
3 Pointwise characterisation of the limit of the derivatives
In this section, we study the convergence of the (directional) derivatives from 2.10 and provide properties for the limits.
Proposition 3.1.
There exists an $\alpha\in H^{1}_{0}(\Omega)$ such that, for a subsequence (that we relabel) the derivative satisfies
$S_{\rho}^{\prime}(f_{\rho})(d)\rightharpoonup\alpha\quad\text{in $H^{1}_{0}(\Omega)$}.$
Proof.
This follows directly from the bound in 2.10. ∎
We want to know what properties the limit $\alpha$ satisfies.
3.1 A complementarity condition on $\alpha$
We start with a simple lemma.
Lemma 3.2.
If $\sigma\colon\mathbb{R}\to\mathbb{R}$ is $C^{1}$ and convex, and satisfies $\sigma=0$ on $(-\infty,r_{0}]$, we have
$\sigma(r)\leq(r-r_{0})\sigma^{\prime}(r)\quad\forall r\in\mathbb{R}.$
Proof.
As $\sigma$ is a differentiable convex function, it satisfies $\sigma(r)\leq\sigma(s)+(r-s)\sigma^{\prime}(r)$ for all $r,s\in\mathbb{R}$. The result follows from the fact that $\sigma(r_{0})=0$. ∎
In the smooth setting where $\lambda_{\rho}(x,\cdot)$ is $C^{1}$, we can take in 3.2 $\sigma(\cdot)=\lambda_{\rho}(x,\cdot)$ for fixed $x$ and obtain
$\lambda_{\rho}(x,r)\leq(r-k_{\rho}(x,\rho))\lambda_{\rho}^{\prime}(x,r)\quad\forall r\in[k_{\rho}(x),K_{\rho}(x)].$ (32)
Note that this inequality also holds in the nonsmooth case, since this implies $k_{\rho}\equiv K_{\rho}$.
Proposition 3.3.
Let $f_{\rho}\to f$ in $H^{-1}(\Omega)$. With $\xi:=f+\Delta u$ and $\alpha$ denoting the weak limit of (a subsequence of) $S_{\rho}^{\prime}(f_{\rho})(d)$ from 3.1, we have
$\langle\xi,\alpha\rangle=\langle\xi,\alpha^{+}\rangle=\langle\xi,\alpha^{-}\rangle=0.$
For a characterisation of this result in terms of the strictly active set, see (39).
Proof.
Set $u_{\rho}:=S_{\rho}(f_{\rho})$ and define the sets
$M_{0}=\{u_{\rho}-\psi\leq k_{\rho}\},\quad M_{1}=\{k_{\rho}<u_{\rho}-\psi\leq K_{\rho}\},\quad M_{2}=\{K_{\rho}\leq u_{\rho}-\psi\}.$
Note that $M_{1}=\emptyset$ in the nonsmooth case. Further, (11) implies
[…]
Define $\xi_{\rho}:=f_{\rho}+\Delta u_{\rho}$. We have
$\langle\xi_{\rho},\alpha_{\rho}\rangle$ $=\int_{\Omega}\Lambda_{\rho}(u_{\rho}-\psi)\alpha_{\rho}\mathop{}\!\mathrm{d}x$
$=\int_{M_{1}}\Lambda_{\rho}(u_{\rho}-\psi)\alpha_{\rho}\mathop{}\!\mathrm{d}x+\int_{M_{2}}\frac{u_{\rho}-\psi+j_{\rho}}{\rho}\alpha_{\rho}\mathop{}\!\mathrm{d}x$
$=\int_{M_{1}}\Lambda_{\rho}(u_{\rho}-\psi)\alpha_{\rho}\mathop{}\!\mathrm{d}x+\int_{M_{2}}\frac{(u_{\rho}-\psi+j_{\rho})^{+}}{\rho}\alpha_{\rho}\mathop{}\!\mathrm{d}x$
with no integral over $M_{0}$ because $\lambda_{\rho}(x,\cdot)$ vanishes on $(-\infty,k_{\rho}(x)]$ and the positive part is introduced due to (33).
Regarding the first term, using the estimate (32), we have
$\left|\int_{M_{1}}\Lambda_{\rho}(u_{\rho}-\psi)\alpha_{\rho}\right|\mathop{}\!\mathrm{d}x$ $\leq\int_{M_{1}}\lambda_{\rho}(\cdot,u_{\rho}-\psi)|\alpha_{\rho}|\mathop{}\!\mathrm{d}x$
$\leq\int_{M_{1}}(u_{\rho}-\psi-k_{\rho})\lambda^{\prime}_{\rho}(\cdot,u_{\rho}-\psi)|\alpha_{\rho}|\mathop{}\!\mathrm{d}x$
$\leq\frac{\left\lVert K_{\rho}-k_{\rho}\right\rVert_{L^{2}(\Omega)}}{\sqrt{\rho}}\left\lVert\sqrt{\lambda_{\rho}^{\prime}(\cdot,u_{\rho}-\psi)}\alpha_{\rho}\chi_{M_{1}}\right\rVert_{L^{2}(\Omega)}$
where we used the bounds on $\lambda_{\rho}^{\prime}$ given in (16) to justify the last estimate. As for the second integral, we obtain
$\left|\int_{M_{2}}\frac{(u_{\rho}-\psi+j_{\rho})^{+}}{\rho}\alpha_{\rho}\mathop{}\!\mathrm{d}x\right|$ $\leq\frac{\left\lVert(u_{\rho}-\psi+j_{\rho})^{+}\chi_{M_{2}}\right\rVert_{L^{2}(\Omega)}}{\sqrt{\rho}}\frac{\left\lVert\alpha_{\rho}\chi_{M_{2}}\right\rVert_{L^{2}(\Omega)}}{\sqrt{\rho}}.$
Thus
$|\langle\xi_{\rho},\alpha_{\rho}\rangle|$ $\leq\frac{\left\lVert K_{\rho}-k_{\rho}\right\rVert_{L^{2}(\Omega)}}{\sqrt{\rho}}\left\lVert\sqrt{\lambda_{\rho}^{\prime}(\cdot,u_{\rho}-\psi)}\alpha_{\rho}\chi_{M_{1}}\right\rVert_{L^{2}(\Omega)}$
$\quad+\frac{\left\lVert(u_{\rho}-\psi+j_{\rho})^{+}\chi_{M_{2}}\right\rVert_{L^{2}(\Omega)}}{\sqrt{\rho}}\frac{\left\lVert\alpha_{\rho}\chi_{M_{2}}\right\rVert_{L^{2}(\Omega)}}{\sqrt{\rho}}.$ (34)
By (10), the first multiplicand in the first term above vanishes as $\rho\to 0$. Let us show that the first multiplicand in the second term also vanishes. By using the triangle inequality and the fact that $(a+b)^{+}\leq a^{+}+b^{+}$ for all $a,b\in\mathbb{R}$,
$\frac{\left\lVert(u_{\rho}-\psi+j_{\rho})^{+}\chi_{M_{2}}\right\rVert_{L^{2}(\Omega)}}{\sqrt{\rho}}$ $\leq\frac{\left\lVert(u_{\rho}-\psi)^{+}\right\rVert_{L^{2}(\Omega)}}{\sqrt{\rho}}+\frac{\left\lVert j_{\rho}\right\rVert_{L^{2}(\Omega)}}{\sqrt{\rho}}$
and both terms on the right-hand side vanish by 2.15 and (20) respectively.
Now we simply need to show that second multiplicand in the two terms on the right-hand side of (34) are bounded. Since $\alpha_{\rho}$ is bounded, so is $d+\Delta\alpha_{\rho}$. Hence,
$C$ $\geq|\langle d+\Delta\alpha_{\rho},\alpha_{\rho}\rangle|=\int_{\Omega}\Lambda_{\rho}^{\prime}(u_{\rho}-\psi)(\alpha_{\rho})\alpha_{\rho}\mathop{}\!\mathrm{d}x$
$\geq\int_{M_{1}}\lambda_{\rho}^{\prime}(\cdot,u_{\rho}-\psi)(\alpha_{\rho})\alpha_{\rho}\mathop{}\!\mathrm{d}x+\int_{M_{2}}\lambda_{\rho}^{\prime}(\cdot,u_{\rho}-\psi)(\alpha_{\rho})\alpha_{\rho}\mathop{}\!\mathrm{d}x,$
where we used that the integral over $M_{0}$ is nonnegative due to (17). Now, on $M_{1}$, the derivative is linear w.r.t. the direction and on $M_{2}$, the derivative is $1/\rho$. Thus, we continue with
$C$ $\geq\int_{M_{1}}\lambda_{\rho}^{\prime}(\cdot,u_{\rho}-\psi)\alpha_{\rho}^{2}+\frac{1}{\rho}\int_{M_{2}}\alpha_{\rho}^{2}\mathop{}\!\mathrm{d}x$
$=\left\lVert\sqrt{\lambda_{\rho}^{\prime}(\cdot,u_{\rho}-\psi)}\alpha_{\rho}\chi_{M_{1}}\right\rVert_{L^{2}(\Omega)}^{2}+\frac{1}{\rho}\left\lVert\alpha_{\rho}\chi_{M_{2}}\right\rVert_{L^{2}(\Omega)}^{2}.$
Plugging this information back into (34), we find that $\langle\xi_{\rho},\alpha_{\rho}\rangle\to 0$ and hence, since $\alpha_{\rho}\rightharpoonup\alpha$ in $H^{1}_{0}(\Omega)$ and $\xi_{\rho}\to\xi$ in $H^{-1}(\Omega)$,
$\langle\xi,\alpha\rangle=0.$
In order to produce the same identity for $\alpha^{+}$, one can repeat exactly the same calculation with $\alpha_{\rho}^{+}$ instead of $\alpha_{\rho}$. In the final step, we can use that $\alpha_{\rho}\rightharpoonup\alpha$ in $H_{0}^{1}(\Omega)$ implies $\alpha_{\rho}^{+}\rightharpoonup\alpha^{+}$ in $H_{0}^{1}(\Omega)$. This yields the claim. ∎
3.2 An orthogonality condition on $-\Delta\alpha-d$
In this section we will show that $-\Delta\alpha-d$ satisfies an orthogonality condition involving the coincidence set.
Lemma 3.4.
We have
$\langle-\Delta\alpha_{\rho}-d,v\rangle=0\quad\forall v\in H^{1}_{0}(\Omega),\;\;v=0\text{ q.e. on $\{(u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)})^{-}=0\}$}.$
Note that the set $\{(u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)})^{-}=0\}$ is defined up to sets of capacity zero, since $u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)}\in H^{1}(\Omega)$ is quasicontinuous, see [MR2731611, Theorem 6.1, §8].
Proof.
Taking $v\in H^{1}_{0}(\Omega)$ such that $v=0$ q.e. on $\{(u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)})^{-}=0\}$, we have
$\langle-\Delta\alpha_{\rho}-d,v\rangle=-\int_{\{u_{\rho}-\psi\geq k_{\rho}\}}\Lambda_{\rho}^{\prime}(u_{\rho}-\psi)(\alpha_{\rho})v\mathop{}\!\mathrm{d}x$ (35)
because $\Lambda_{\rho}^{\prime}(u_{\rho}-\psi)$ vanishes when $u_{\rho}-\psi\leq k_{\rho}$. By definition, we have that $v=0$ q.e. on the set $\{u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)}\geq 0\}$. Note that $u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)}\geq u_{\rho}-\psi-k_{\rho}$ a.e. and so we have the inclusion
$\{u_{\rho}-\psi-k_{\rho}\geq 0\}\subset\{u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)}\geq 0\}.$
It follows that $v=0$ a.e. on $\{u_{\rho}-\psi-k_{\rho}\geq 0\}$ too. Using this in (35) we obtain the desired statement. ∎
Proposition 3.5.
We have
$\langle-\Delta\alpha-d,v\rangle$ $=0\qquad\text{$\forall v\in H^{1}_{0}(\Omega):v=0$ q.e. on $\{u=\psi\}$}.$
Proof.
By (9) it follows that $u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)}\to u-\psi$ in $H^{1}(\Omega)$. We want to show that $s_{\rho}:=(u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)})^{-}$ converges to $s:=(u-\psi)^{-}=\psi-u$ in capacity.
From 2.14 and [TowardsM, Lemma 2.2], we get that $u_{\rho}$ converges towards $u$ in capacity, i.e.,
$\operatorname*{cap}(\{|u_{\rho}-u|\geq\varepsilon\})\to 0\quad\text{as }\rho\to 0$
for all $\varepsilon>0$. Let $\varepsilon>0$ be arbitrary. By (9), there exists $\rho_{0}>0$ such that $\rho\leq\rho_{0}$ implies $\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)}\leq\frac{\varepsilon}{2}$. Restricting to such sufficiently small $\rho$, we have
$\lvert s_{\rho}-s\rvert\leq\lvert u_{\rho}+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)}-u\rvert\leq\lvert u_{\rho}-u\rvert+\frac{\varepsilon}{2}.$
Consequently,
$\operatorname*{cap}(\{|s_{\rho}-s|\geq\varepsilon\})\leq\operatorname*{cap}(\{|u_{\rho}-u|\geq\varepsilon/2\})\to 0\quad\text{as }\rho\to 0.$
So we have shown that $s_{\rho}\to s$ in capacity as desired.
Now, the result of 3.4 implies
±$\pm(-\Delta\alpha_{\rho}-d)\in\{v\in H^{1}_{0}(\Omega):v\geq 0\text{ q.e. on $\Omega$ and }v=0\text{ q.e. on }\{s_{\rho}=0\}\}^{\circ}$
with $s_{\rho}=(u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)})^{-}$ as above. Applying [TowardsM, Lemma 2.6] we find
±$\pm(-\Delta\alpha-d)\in\{v\in H^{1}_{0}(\Omega):v\geq 0\text{ q.e. on $\Omega$and }v=0\text{ q.e. on }\{s=0\}\}^{\circ}$
where $s=(u-\psi)^{-}=\psi-u$ is the limit (w.r.t. convergence in capacity) of $s_{\rho}$. This yields the desired statement after using the decomposition $v=v^{+}-v^{-}$. ∎
Remark 3.6.
The proof of 3.4 even shows
$\langle-\Delta\alpha_{\rho}-d,v\rangle=0\quad\forall v\in H^{1}_{0}(\Omega),\;\;v=0\text{ a.e. on $\{(u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}(\Omega)})^{-}=0\}$}.$
However, one cannot generalise (the proof of) 3.5 to obtain
$\langle-\Delta\alpha-d,v\rangle$ $=0\qquad\text{$\forall v\in H^{1}_{0}(\Omega):v=0$ a.e. on $\{u=\psi\}$}.$
The problem is the following. The set $\{u=\psi\}$ could have measure zero but positive capacity. In this case, the assumption $v=0$ a.e. on $\{u=\psi\}$ is void. Consequently, $v$ could be constant $1$ in a neighbourhood of $\{u=\psi\}$. If further $v_{\rho}\to v$ in $H_{0}^{1}(\Omega)$ is given (this is needed for passing to the limit with $\langle-\Delta\alpha_{\rho}-d,v\rangle$), we have $v_{\rho}\to v$ in capacity. Further, $\{u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}}\geq 0\}$ could contain an open neighbourhood of $\{u=\psi\}$. Therefore, $v_{\rho}=0$ a.e. on $\{u_{\rho}-\psi+\left\lVert k_{\rho}\right\rVert_{L^{\infty}}\geq 0\}$ implies $v_{\rho}=0$ q.e. on $\{u=\psi\}$. This contradicts $v_{\rho}\to v$ in capacity.
If $\psi\in H^{1}_{0}(\Omega)$, the argument above can be simplified and we could also alternatively have argued in a similar way to [AHRW, §7].
4 Generalised derivatives as limits
In the convergence results of the previous section, the direction $d$ was fixed. We need something stronger than this: we would like limiting statements for the derivative when seen as a linear operator, and furthermore, we would like the subsequence that converges to be independent of $d$.
4.1 Existence of limiting elements
First, we prove that the Gâteaux derivatives converge in the weak operator topology.
Proposition 4.1.
Let a sequence $\{f_{\rho}\}\subset H^{-1}(\Omega)$ be given such that $f_{\rho}\to f$ in $H^{-1}(\Omega)$ and such that $S_{\rho}$ is Gâteaux differentiable at $f_{\rho}$ for all $\rho>0$. Then, there exists a map $L\in\mathcal{L}(H^{-1}(\Omega),H^{1}_{0}(\Omega))$ such that for a subsequence (that we relabel),
$S_{\rho}^{\prime}(f_{\rho})$ $\overset{\mathrm{WOT}}{\to}L.$
Proof.
The bound of 2.10 shows that
$\left\lVert S_{\rho}^{\prime}(f_{\rho})\right\rVert_{\mathcal{L}(H^{-1}(\Omega),H^{1}_{0}(\Omega))}\leq 1$
for all $\rho>0$. By the separability of $H^{1}_{0}(\Omega)$, one can show that the unit ball is sequentially compact in the WOT, see also [CCMW, Remark 3.1]. The result follows. ∎
Note that the Gâteaux differentiability assumption of the previous result holds automatically in the smooth case; this follows from (21). The question remains whether it can also hold in the nonsmooth case. We can in fact use the generalised Rademacher’s theorem of Mignot [Mignot, Theorem 1.2], which tells us that for each $\rho$, there exists a dense set on which $S_{\rho}$ is Gâteaux differentiable. However, this dense set depends on $\rho$, and because we wish to consider limits it would be ideal to have one dense set on which $S_{\rho}$ is Gâteaux for all $\rho$ (or at least for a sequence $\{\rho_{n}\}$). We cannot simply take the intersection because an intersection of dense sets may not be dense. Thus we need another argument where we essentially embed $\{S_{\rho}\}$ into an infinite dimensional space, which we give next.
Lemma 4.2.
Let a sequence $\{\rho_{n}\}\subset(0,1]$ with $\rho_{n}\to 0$ be given and, for brevity, set $S_{n}:=S_{\rho_{n}}$. Then, there exists a dense set $F\subset H^{-1}(\Omega)$ such that $S_{n}$ is Gâteaux differentiable on $F$ for all $n\in\mathbb{N}$.
We remark that this result also follows from the theory of Aronszajn null sets. Indeed, because $S_{\rho}$ is Lipschitz, $S_{\rho}$ is Gâteaux differentiable outside an Aronszajn null set [MR1727673, Theorem 6.42] and the countable union of Aronszajn null sets remains Aronszajn null (see the paragraph after Definition 6.23 of [MR1727673]). This implies the result. For convenience, however, we give a different and direct proof.
Proof.
We denote by $\mathcal{V}:=\ell^{2}(\mathbb{N};H^{1}_{0}(\Omega))$ the set of $\ell^{2}$-summable $H^{1}_{0}(\Omega)$-valued sequences. Consider the map $H\colon H^{-1}(\Omega)\to\mathcal{V}$ defined by
$H(f):=\big{(}2^{-1/2}S_{1}(f),2^{-2/2}S_{2}(f),\ldots\big{)},$
i.e., the $n$th component is $[H(f)]_{n}=2^{-n/2}S_{n}(f)$. Note that
$\left\lVert H(f)\right\rVert_{\mathcal{V}}^{2}=\sum_{n=1}^{\infty}\left\lVert 2^{-n/2}S_{n}(f)\right\rVert_{H^{1}_{0}(\Omega)}^{2}\leq C^{2}\sum_{n=1}^{\infty}2^{-n}=C^{2}<\infty,$
using the fact that $\left\lVert S_{n}(f)\right\rVert_{H^{1}_{0}(\Omega)}\leq C$ for a constant $C$ independent of $n$, see 2.9. Further, $H$ is Lipschitz, since
$\left\lVert H(f)-H(g)\right\rVert_{\mathcal{V}}^{2}\leq\sum_{n=1}^{\infty}2^{-n}\left\lVert(S_{n}(f)-S_{n}(g))\right\rVert_{H^{1}_{0}(\Omega)}^{2}\leq\left\lVert f-g\right\rVert_{H^{-1}(\Omega)}^{2},$
where we used the fact that each $S_{n}$ is Lipschitz with constant $1$, see 2.9. Since $\mathcal{V}$ is a Hilbert space and $H^{-1}(\Omega)$ is a separable Hilbert space, it follows [Mignot, Theorem 1.2] that $H$ is Gâteaux differentiable on a dense set $F\subset H^{-1}(\Omega)$. Now, it is straightforward to check that the Gâteaux differentiability of $H$ on $F$ implies that $S_{n}$ is Gâteaux differentiable on $F$ for all $n\in\mathbb{N}$. ∎
Observe that this result holds for a countable sequence and does not and cannot hold for the original uncountable family $\{S_{\rho}\}$. This is because the differentiability holds up to a set of measure zero and countable unions (of the exceptional sets) still have measure zero, which of course, does not apply for the uncountable case. A simple real-valued example is $S_{\rho}(x):=|x-\rho|.$ This function is not differentiable at $\rho$, so if one considers all of these functions, the common differentiability set is just the empty set.
4.2 Capacitary measures and characterisation of limits
In what follows, we assume that the reader is familiar with the notions of capacity, quasi-open and quasi-closed sets and quasicontinuous representatives. For an introduction tailored to optimal control, we refer to [HarderWachsmuth, Section 3].
The measure $\mu_{\rho}$ defined in (25) has some additional regularity that we can exploit. We begin with recalling the notion of capacitary measures, see, e.g., [RW, Definition 3.1].
Definition 4.3 (Capacitary measure).
A capacitary measure is a Borel measure $\mu$ such that
  1. (i)
    $\mu(B)=0$ for every Borel set $B\subset\Omega$ with $\operatorname*{cap}(B)=0$
  2. (ii)
    $\mu$ is regular in the sense that every Borel set $B\subset\Omega$ satisfies
    $\mu(B)=\inf\{\mu(O):O\text{ quasi-open and }\operatorname*{cap}(B\setminus O)=0\}.$
We denote by $\mathcal{M}_{0}(\Omega)$ the set of all capacitary measures on $\Omega$.
Note that a measure which is absolutely continuous w.r.t. the Lebesgue measure is a capacitary measure, see also [DalMaso, Remark 3.2 and Definition 2.1]. In particular, (25) defines a capacitary measure in the setting of 2.11.
From now on, given (an equivalence class) $v\in H_{0}^{1}(\Omega)$, we will always work with a Borel measurable and quasicontinuous representative. Note that such a representative is uniquely determined up to subsets of capacity zero which are $\mu$-nullsets for every $\mu\in\mathcal{M}_{0}(\Omega)$.
We denote by $L^{2}_{\mu}(\Omega)$ the usual Lebesgue space w.r.t. a measure $\mu\in\mathcal{M}_{0}(\Omega)$.
For $\mu\in\mathcal{M}_{0}(\Omega)$ and $f\in H^{-1}(\Omega)$, one can check that there is a unique solution $y\in H_{0}^{1}(\Omega)\cap L^{2}_{\mu}(\Omega)$ of
$\langle-\Delta y,v\rangle+\int_{\Omega}yv\mathop{}\!\mathrm{d}\mu=\langle f,v\rangle\qquad\forall v\in H_{0}^{1}(\Omega)\cap L^{2}_{\mu}(\Omega).$ (36)
The solution map of this weak formulation is denoted by $L_{\mu}\colon H^{-1}(\Omega)\to H^{1}_{0}(\Omega)$, i.e., $L_{\mu}f=y$.
Associated to the VI (1), define the inactive set (or non-coincidence set) $I:=\{u<\psi\}$ and the active set (or coincidence set) $A:=\{u=\psi\}$. We denote by $A_{s}\subset A$ the strictly active set which can be defined via the quasi-support (or fine support) of the measure $\xi=f+\Delta u$; this is a quasi-closed subset. Note that all these sets are defined up to subsets of capacity zero.
Under 1.2 and if $S(f)\in C_{0}(\Omega)$, the strong-weak generalised derivative $\partial_{B}^{sw}S(f)\subset\mathcal{L}(H^{-1}(\Omega),H^{1}_{0}(\Omega))$ of $S$ at $f$ was characterised in [RW, Theorem 5.6] and we have
$\partial_{B}^{sw}S(f)=\{L_{\mu}:\mu\in\mathcal{M}_{0}(\Omega),\;\mu(I)=0,\;\mu=\infty\text{ on }A_{s}\},$ (37)
where $\mu=\infty$ on $A_{s}$ is to be understood in the sense that
$v=0\quad\text{q.e. on $A_{s}$ for all $v\in H^{1}_{0}(\Omega)\cap L^{2}_{\mu}(\Omega)$}$
(see [RW, Lemma 5.2] for some equivalent characterisations). We will show that the limits of the derivatives from 4.1 belong to this set.
In the linear setting of 2.11, $\mu_{\rho}$ is a measure and the equation (26) for $\alpha_{\rho}$ is to be understood in the weak form
$\langle-\Delta\alpha_{\rho},z\rangle+\int_{\Omega}\alpha_{\rho}z\mathop{}\!\mathrm{d}\mu_{\rho}=\langle d,z\rangle\qquad\forall z\in H^{1}_{0}(\Omega).$
Note that (25) implies
$\int_{\Omega}\alpha_{\rho}z\mathop{}\!\mathrm{d}\mu_{\rho}=\int_{\Omega}\Lambda_{\rho}^{\prime}(u_{\rho}-\psi)\alpha_{\rho}z\mathop{}\!\mathrm{d}x.$
Since $\Lambda_{\rho}^{\prime}(u_{\rho}-\psi)$ belongs to $L^{\infty}(\Omega)$, see 2.11, we have $H_{0}^{1}(\Omega)\subset L^{2}_{\mu_{\rho}}(\Omega)$. Consequently, the above equation is equivalent to $L_{\mu_{\rho}}(d)=\alpha_{\rho}$ and this implies
$L_{\mu_{\rho}}\equiv S_{\rho}^{\prime}(f_{\rho}).$ (38)
To summarise, if the directional derivative $S_{\rho}^{\prime}(f_{\rho})$ is linear, we can find a capacitary measure $\mu_{\rho}$ such that (38) holds.
We now define a notion of convergence for measures related to convergence in the weak operator topology (see 1.1) of the associated solution operators $L_{\mu}$ defined in (36).
Definition 4.4.
[…]
The sequential compactness of the space $\mathcal{M}_{0}(\Omega)$, see, e.g., [DalMasoMosco, Theorem 4.14], is a crucial property which implies the next result.
Proposition 4.5.
[…]
The key observation that allows us to prove the next theorem, which is the main result, is the following. 4.5 implies, under the stated assumptions, that there is a subsequence (which we shall relabel) such that for every $d\in H^{-1}(\Omega)$, $\alpha_{\rho}:=S_{\rho}^{\prime}(f_{\rho})d\rightharpoonup L_{\mu}d=:\alpha$. We emphasise that the subsequence is independent of $d$. Thus, all of the results of Section 3 can be applied to obtain information on $\alpha$ and, consequently, on the limiting operator $L_{\mu}$.
We come now to our main result, which concatenates the above results and characterises the limiting elements obtained in 4.1. We will use the fact that the result of 3.3 is equivalent [HarderWachsmuth, Lemma 3.7] to
[…]
Theorem 4.6.
Let $f_{\rho}\to f$ in $H^{-1}(\Omega)$ with $S(f)\in C_{0}(\Omega)$ be given. We further assume that $S_{\rho}$ is Gâteaux differentiable at $f_{\rho}$ for all $\rho>0$. Then there exists a map $L\in\partial_{B}^{sw}S(f)$ such that, for a subsequence (that we relabel),
$S_{\rho}^{\prime}(f_{\rho})\overset{\mathrm{WOT}}{\to}L.$
Proof.
Recalling the characterisation
$\partial_{B}^{sw}S(f)=\{L_{\mu}:\mu\in\mathcal{M}_{0}(\Omega),\;\mu(I)=0,\;\mu=\infty\text{ on }A_{s}\},$
we will show that the operator $L_{\mu}$ from 4.5 belongs to the set on the right-hand side.
We choose a nonnegative $v\in H_{0}^{1}(\Omega)$ such that $\{v>0\}=I$. Further, set $d:=-\Delta v$ and $\alpha:=L_{\mu}d$, i.e.,
$-\Delta\alpha+\mu\alpha=-\Delta v.$
Testing with $\alpha$ we get $\left\lVert\nabla\alpha\right\rVert_{L^{2}}\leq\left\lVert\nabla v\right\rVert_{L^{2}}$. Consequently,
$\langle-\Delta(\alpha-v),\alpha-v\rangle\leq 2\langle-\Delta v,v\rangle+2\langle\Delta\alpha,v\rangle=2\langle d,v\rangle+2\langle\Delta\alpha,v\rangle=0$
with the final equality by 3.5. Thus, $\alpha=v$. Testing the equation for $\alpha$ by $v$, we get
$\int_{\Omega}v^{2}\mathop{}\!\mathrm{d}\mu=0$
and this means $\mu(I)=0$.
Now, we choose $d:=1$ and $\alpha:=L_{\mu}d$. Note that this $\alpha$ coincides with $w_{\mu}:=L_{\mu}(1)$ defined below Theorem 3.8 of [RW]. From the characterisation (39) of the result of 3.3, we get $\alpha=w_{\mu}=0$ q.e. on $A_{s}$ and by [RW, Lemma 5.2] we get $\mu=+\infty$ on $A_{s}$. ∎
In the nonsmooth case, Theorem 4.6 and 4.2 imply that there exists a dense subset $F\subset H^{-1}(\Omega)$ such that if $f\in F$ and $S(f)\in C_{0}(\Omega)$, there exists a map $L\in\partial_{B}^{sw}S(f)$ such that, for a subsequence (that we relabel),
$S_{n}^{\prime}(f)$ $\overset{\mathrm{WOT}}{\to}L.$
5 Optimal control of the obstacle problem
We apply our findings to the optimal control of the obstacle problem. Let $F_{ad}\subset L^{2}(\Omega)$ be a nonempty, closed and convex set satisfying
$S(f)\in C_{0}(\Omega)\qquad\forall f\in F_{ad}$
(see Remark 1.4) and let $J\colon H^{1}_{0}(\Omega)\times L^{2}(\Omega)\to\mathbb{R}$ be a given objective function satisfying
  1. (i)
    $J$ is continuously Fréchet differentiable with partial derivatives $J_{y}$ and $J_{f}$,
  2. (ii)
    If $y_{n}\to y$ in $H^{1}_{0}(\Omega)$ and $f_{n}\rightharpoonup f$ in $L^{2}(\Omega)$, then
    $J(y,f)\leq\liminf_{n\to\infty}J(y_{n},f_{n}).$
An example of $J$ satisfying the above conditions is
$J(y,f):=\frac{1}{2}\left\lVert y-y_{d}\right\rVert_{L^{2}(\Omega)}^{2}+\frac{\nu}{2}\left\lVert f\right\rVert_{L^{2}(\Omega)}^{2},$
where $y_{d}\in L^{2}(\Omega)$ is a given desired state and $\nu\geq 0$ is a constant.
Consider the optimal control problem
$\min_{f\in F_{ad}}J(y,f)\quad\text{such that}\quad y=S(f).$ (40)
In the next result, we will derive a first-order optimality condition for this control problem. This rigorously shows the satisfaction of [RW, (19)], which was derived only formally there.
Theorem 5.1.
For any local minimiser $(\bar{y},\bar{f})\in H^{1}_{0}(\Omega)\times F_{ad}$ of (40), there exists $L\in\partial_{B}^{sw}S(\bar{f})$ such that
$0\in L^{*}J_{y}(\bar{y},\bar{f})+J_{f}(\bar{y},\bar{f})+N_{F_{ad}}(\bar{f})$
is satisfied where $N_{F_{ad}}(\bar{f})$ is the normal cone to $F_{ad}$ at $\bar{f}$.
Proof.
We denote by $\varepsilon>0$ the radius of optimality, i.e.,
$J(S(f),f)\geq J(\bar{y},\bar{f})\qquad\forall f\in F_{ad}\cap B_{\varepsilon}(\bar{f}),$
where $B_{\varepsilon}(\bar{f})$ is a closed ball in $L^{2}(\Omega)$.
We regularise the problem (40) and consider
$\min_{f\in F_{ad}\cap B_{\varepsilon}(\bar{f})}J(S_{\rho}(f),f)+\frac{1}{2}\left\lVert f-\bar{f}\right\rVert_{L^{2}(\Omega)}^{2},$
where we take $S_{\rho}$ to be $S_{\rho}^{\mathrm{sm}}$ or $S_{\rho}^{\mathrm{sc}}$. Denote by $f_{\rho}$ a global minimiser of this problem and set $y_{\rho}:=S_{\rho}(f_{\rho})$.
Using standard arguments, one can show $f_{\rho}\to\bar{f}$ in $L^{2}(\Omega)$. Consequently, the constraint $f_{\rho}\in B_{\varepsilon}(\bar{f})$ is not binding for small enough $\rho$. By the standard minimisation principle, we get
$0\in\bar{J}^{\prime}(f_{\rho})+(f_{\rho}-\bar{f})+N_{F_{ad}}(f_{\rho})$
where $\bar{J}(f):=J(S_{\rho}(f),f)$ defines the reduced functional. Using the chain rule and the fact that $J$ is Fréchet, we can write this as
$0\in S_{\rho}^{\prime}(f_{\rho})^{*}J_{y}(y_{\rho},f_{\rho})+J_{f}(y_{\rho},f_{\rho})+(f_{\rho}-\bar{f})+N_{F_{ad}}(f_{\rho})\quad\text{in }L^{2}(\Omega).$
Note that $J_{y}(y_{\rho},f_{\rho})\in H^{-1}(\Omega)$ and $S_{\rho}^{\prime}(f_{\rho})^{*}\colon H^{-1}(\Omega)\to H_{0}^{1}(\Omega)\subset L^{2}(\Omega)$.
It remains to pass to the limit with this optimality condition. From 2.14 we get $y_{\rho}\to\bar{y}$ in $H_{0}^{1}(\Omega)$. Further, Theorem 4.6 enables us to select a subsequence (which we relabel) such that $S_{\rho}^{\prime}(f_{\rho})\overset{\mathrm{WOT}}{\to}L$ for some $L\in\partial_{B}^{sw}S(\bar{f})$. Further, we note that $S_{\rho}^{\prime}(f_{\rho})$ and $L$ are self-adjoint, which yields $S_{\rho}^{\prime}(f_{\rho})^{*}\overset{\mathrm{WOT}}{\to}L^{*}$ Together with the product rule [RW, Lemma 2.9 (ii)] we get
$S_{\rho}^{\prime}(f_{\rho})^{*}J_{y}(y_{\rho},f_{\rho})+J_{f}(y_{\rho},f_{\rho})+(f_{\rho}-\bar{f})\rightharpoonup L^{*}J_{y}(\bar{y},\bar{f})+J_{f}(\bar{y},\bar{f})$
in $H_{0}^{1}(\Omega)$ and, consequently, strongly in $L^{2}(\Omega)$. Since the graph of the normal cone map is closed, this implies
$0\in L^{*}J_{y}(\bar{y},\bar{f})+J_{f}(\bar{y},\bar{f})+N_{F_{ad}}(\bar{f})$
as claimed. ∎
Thanks to this result, [RW, Lemma 7.2] implies the existence of
$p\in H^{1}_{0}(\Omega\setminus A_{s}),\qquad\nu\in H^{-1}(\Omega),\qquad\lambda\in N_{F_{ad}}(\bar{f}),$
such that
$p+J_{f}(\bar{y},\bar{f})+\lambda$ $=0,$ $\langle\nu,v\rangle$ $=0\quad\forall v\in H^{1}_{0}(\Omega\setminus A),$
$J_{y}(\bar{y},\bar{f})+\Delta p-\nu$ $=0,$ $\langle\nu,p\varphi\rangle$ $\geq 0\quad\forall\varphi\in W^{1,\infty}(\Omega)^{+},$
which is a necessary condition satisfied by every local minimiser as shown in [SchielaWachsmuth]. Here, $A_{s}$ and $A$ denote the strictly active set and the active set associated with $(\bar{y},\bar{f})$, respectively.
[…]